### v1

# Reducing Size Bias in Sampling for Infectious Disease Spread on Networks

Abstract
Epidemiological models can aid policymakers in reducing disease spread by predicting outcomes based on disease dynamics and contact network characteristics. Accurate calibration of these models requires representative samples of contact networks. In this connection, we investigate the performance of two sampling algorithms—Random Walk (RW) and Metropolis-Hastings Random Walk (MHRW)—across three network types: Erdős–Rényi (ER), Small-World (SW), and Scale-Free (SF). Disease transmission is simulated using a stochastic susceptible-infected-recovered (SIR) framework. Our findings show that for ER and SW networks, RW overestimates infected individuals and secondary infections by approximately $25\%$, while MHRW produces estimates that align well with the underlying network structure. Both methods also yield estimates closer to the underlying network for time-to-infection. However, sampling SF networks introduces significant variability for both algorithms. Removing duplicate samples further reduces MHRW’s accuracy in all network types. We apply both algorithms to a cattle movement network consisting of 46,512 farms, exhibiting characteristics of ER, SW, and SF networks. RW overestimates infected farms by approximately $100\%$ and secondary infections by $>900\%$, whereas MHRW reduces these overestimations to approximately $50\%$ and $<800\%$, respectively. Time-to-infection estimates reveal that RW underestimates by approximately $40\%$, while MHRW slightly overestimates by $10\%$. These results highlight the importance of selecting sampling algorithms based on network structure and disease severity. RW provides conservative estimates suited for high-mortality, fast-spreading diseases, while MHRW is more suitable for precise interventions in less severe outbreaks. These insights can guide policymakers in optimising resource allocation and disease control strategies.
keywords:
networks , sampling , disease modelling , size bias , cattle network , policymaking , SIR model
MSC:
92D30 , 05C82 , 62P10 , 91D30 , 90B15
[…]
[inst1]organization=School of Mathematics, Cardiff Univerity,addressline=Senghennydd Road, city=Cardiff, postcode=CF24 4AG, country=UK
1 Introduction
During an epidemic, accurately estimating metrics such as the number of infected individuals, effective reproduction number, and time to infection is crucial for preparedness (Nunes et al., 2024). Mathematical modelling provides a suite of tools, frameworks, and methodologies to generate these estimations (Vynnycky and White, 2010). However, estimating disease metrics on networks is challenging due to factors such as network complexity (Avraam and Hadjichrysanthou, 2025) and the influence of the chosen data sampling method (Danon et al., 2011). In particular, the Random Walk (RW) sampling algorithm is prone to size bias, leading to over-representation of highly connected individuals and potentially skewed estimates of disease spread. In this paper, we statistically compare two sampling algorithms, Random Walk (RW) and Metropolis-Hastings Random Walk (MHRW), using stochastic susceptible-infected-recovered (SIR) model simulations to evaluate the estimation of disease metrics on three network types (Erdős-Rényi (ER), Small-world (SW), and Scale-free (SF)) and discuss the implications for policy planning.
The extent and intensity of disease spread are driven broadly by three factors: the transmissibility of the virus, the recovery rate of individuals, and the contact network within the population (Nelson and Williams, 2014). Mechanistic models have been developed based on various combinations of these factors to study disease dynamics, predict disease trajectories, and examine the sensitivity of parameters such as transmission rate and recovery rate (Boccaletti et al., 2006, Kiss et al., 2017a, b).
Contact networks act as a means for infectious disease transmission. In a contact network, individuals are represented as nodes and an edge represents interaction between two nodes. The number of edges (connections) attached to a node is its degree (Newman, 2018). In this work, we use undirected contact networks, i.e., disease transmission can occur in both directions along an edge.
We consider three types of contact networks: 1) ER networks (Erdős and Rényi, 1959), which serve as a simple yet useful benchmark for understanding disease spread in random, unstructured networks; 2) SW networks (Watts and Strogatz, 1998), which closely mimic social mixing networks due to the high clustering and long-range links; and 3) SF networks (Li et al., 2005), characterized by the presence of hubs with high connectivity, reflecting the presence of super-spreaders in a real-world contact network (Lieberthal and Gardner, 2021). SF networks capture the extreme heterogeneity of connections found in real-world contact networks, which is less pronounced in the SW and ER networks, which have relatively uniform connectivity. Although these are theoretical networks they capture many properties of real-world networks (Aldous and Wilson, 2003). Thus, we use them here to demonstrate the diversity of behaviours that can arise.
Since our primary interest is to understand the influence of the underlying contact network on disease transmission metrics, we consider a stochastic SIR model (Brauer, 2008) of disease spread on the contact network. The SIR models are the most commonly used mechanistic compartmental models for infectious disease spread (Kermack and McKendrick, 1927). SIR dynamics can be simulated at either the population level or the individual level; here we do the latter.
A SIR model at the population level consists of a deterministic system of ordinary differential equations (ODEs), which describes the rate of change in the number or proportion of individuals in three compartments: susceptible, infected, and recovered, at the population level. The solution of ODEs predict key features of disease spread, such as the number of infections, the rate of transmission, and the effective reproduction number (Brauer, 2008).
The SIR dynamics at the individual level, as we use in this work, takes the form of stochastic equations describing the rate of change in the probability of the population’s state. The SIR framework at the individual level, accounts for heterogeneous mixing within the population based on the underlying contact network (Kiss et al., 2017a).
Another approach to studying disease spread is to use data to extract correlation patterns using statistical modelling. In statistical models, usually, only a sample of the data is adequate. There are two key reasons for using sample data: a) the large cost and time needed in collecting data for an entire population, and b) the high computational cost of processing large datasets. The accuracy of predictions of both statistical and mechanistic models depends on the quality of the sample; if the sample is not representative of the population then it leads to biased estimates of disease metrics.
Various types of biases can arise if an incorrect sampling method is employed, including demographic bias (Tyrer and Heyman, 2016), geographic bias (Banerjee and Chaudhury, 2010), size bias, and temporal bias (Kleinbaum et al., 1981). For instance, in cohort studies for survival analysis of cancer patients, size bias occurs when individuals with longer survival times are more likely to be included in the sample. This can lead to an overestimation of survival time if the bias is not corrected (Shen et al., 2009).
For diseases which are highly transmissible via close contact such as COVID-19, measles, chickenpox etc., the number of contacts is a critical factor in the spread of the virus, it is essential to obtain an unbiased sample that accurately reflects the population’s contact distribution. Failure to do so can lead to incorrect estimation and prediction of the disease spread rate and ineffective mitigation.
To sample data from a contact network, traversal-based sampling (TBS) methods are widely used (Gjoka et al., 2010, Craft and Caillaud, 2011, Malmros et al., 2016, Fournet and Barrat, 2017, Cui et al., 2022). These methods begin with one or more initial nodes (called seeds) and, based on certain information about their neighbours, choose the next node to sample. The primary distinction among TBS methods lies in how the next node is selected.
Two classic probabilistic TBS methods are the RW (Wei et al., 2004) and MHRW algorithms (Hu and Lau, 2013). The RW and MHRW are Markov chain-based algorithms and share the “memoryless” property, leading to duplicates in the sample. In this work, we also compare the sampling algorithms when we remove or keep the duplicate sample points.
The RW algorithm is often used in epidemiological studies of large populations (Milligan et al., 2004, Qi et al., 2023) as it is relatively fast compared to other sampling algorithms. Specifically, for infectious diseases, the RW algorithm works well for homogeneous contact networks, where individuals have a similar number of contacts. However, real-world contact networks are often heterogeneous (Nielsen et al., 2020), that is, the individuals vary in their number of contacts. This heterogeneity can result in size biased samples when using the RW algorithm, i.e., over-representation of highly connected individuals (high-degree nodes) (Stein et al., 2014, Li et al., 2015).
In (Gjoka et al., 2010), the authors compare the RW and MHRW algorithms for sampling online social networks, specifically Facebook. Their results show that the RW algorithm produces a size biased sample, over-representing high-degree nodes, whereas the MHRW algorithm generates a sample with a degree distribution closely matching that of the underlying Facebook network. Similar observations are reported in (Leskovec and Faloutsos, 2006b) for various network types, including citation networks from the e-print arXiv (Leskovec and Faloutsos, 2006b), autonomous system networks (University of Oregon, ) (Internet router networks), affiliation networks from the astrophysics category on arXiv (Leskovec and Faloutsos, 2006b), and trust networks from Epinions.com (Richardson et al., 2003). We note, however, that similar comparative analyses in the context of epidemiological research are limited. This study seeks to address this shortcoming by exploring the RW and MHRW sampling methods in the domain of infectious disease modelling.
Here, we examine how network structures and sampling algorithms affect estimation of disease metrics. We, thus, simulate a stochastic SIR model with varying transmission and recovery rates for three network types (ER, SW and SF). Specifically, in Section 2, we illustrate the size bias in the measurement of disease metrics introduced by the RW and MHRW algorithms, and in Section 3 we compare key metrics such as epidemic size, the number of secondary infections (effective reproduction number), and the time to infection. Finally, in Section 4, to illustrate the practical implications of our study, we analyse cattle movement network data provided by the British Cattle Movement Service (BCMS), which has hybrid properties linked to the ER, SW and SF networks. We provide a summary and directions for future research in Section 5.
2 Methods
For completeness, we briefly define and describe several commonly used properties of networks. We also describe the RW and MHRW algorithms used for sampling from networks. Additionally, we outline the formulation and simulation methodology of a stochastic SIR model for infectious disease spread on networks.
2.1 Networks
We consider contact networks (Bansal et al., 2010, Craft, 2015). A contact network is defined by $G=\{V,E\}$, where $V$ is a set of node labels representing the individuals who can carry and transmit the disease, and $E$ is a set of edges that defines whether there have been interactions between the individuals. Let $N$ be the number of nodes in $G$, also known as network size. Let $k_{v}$ denote the degree of a node $v\in V$, which is defined as the number of edges (connections) of node $v$. Based on the network structure we can define and calculate (at least numerically) $p_{k}$, the probability that a node is of degree $k$, as well as the average degree of the network $\langle k\rangle$ (Newman, 2018). We will randomly generate our network structures on the basis of these degree distributions. Namely, a number of nodes will be fixed and the connections will be chosen to satisfy a given degree distribution.
2.1.1 Network Structures
Real-world networks, such as the cattle movement network from the BCMS, which we analyze in Section 4, rarely align perfectly with theoretical network structures like ER, SW, or SF networks. Instead, they often have a mixture of features from these networks, such as clustering, and heterogeneous connectivity. To better understand the implications of these structural characteristics on disease spread, we first describe these three theoretical networks individually (Newman, 2018). This approach helps us to determine how specific network features influence disease metrics, offering a baseline for understanding the more complex structure of the cattle movement network and other real networks.
(a) Erdős-Rényi network
(b) Small-world network
(c) Scale-free network
Figure 1: Illustration of three network structures, generated using the NetworkX (Hagberg et al., 2008) library, with $N=50$ nodes and average degree $\langle k\rangle=5$: a) ER network, b) SW network with $p=0.5$, and c) SF network with $\alpha=3$.
(d) Erdős-Rényi network
(e) Small-world network
(f) Scale-free network
Figure 2: Degree distribution (histogram) and fitted theoretical degree distributions (black line plots): a) the binomial distribution given by Eq. (1) with $N=50$ and $p=0.1$ for ER networks, b) a fitted Gaussian distribution with mean $4$ and standard deviation $1.2$, applied to the degree values from generated SW networks, and c) the power-law distribution given by Eq. (2) with $\alpha=3$ for SF networks. Degree distributions are estimated using $100$ simulated networks of each type with parameters $N=50$, $\langle k\rangle=5$ for the ER network, $p=0.5$ for the SW network, and $\alpha=3$ for the SF network, generated using the NetworkX (Hagberg et al., 2008) library.
  1. 1.
    Erdős-Rényi (ER) networks are generated by considering every pair of nodes and connecting them with an edge with probability $p$, independently of other edges in the network (Erdos et al., 1960). An example of an ER network is shown in Figure 1(a). ER networks follow binomial degree distribution (Newman et al., 2001), with network size $N$ and edge creation probability $p=\langle k\rangle/(N-1)$ (Figure 1(d)), that is, given by
    $p_{k}=\binom{N}{k}p^{k}(1-p)^{N-k}.$ (1)
  2. 2.
    Small-world (SW) networks are characterized by high clustering and a small average shortest path distance between nodes (see Figure 1(b)), and they follow a skewed Gaussian degree distribution (Figure 1(e)). Two common methods for generating SW networks are: 1) the Watts-Strogatz (WS) model (Watts and Strogatz, 1998), and 2) the Newman-Watts-Strogatz (NWS) model (Newman and Watts, 1999). In both methods, the process begins with a regular lattice network with average degree $\langle k\rangle$. In the NWS method, edges are added with probability $p$, whereas in the WS method, edges are rewired with probability $p$. For $p=0$, the network is a regular lattice, and for $p=1$, the network is an ER network. For $10<p<$ 1, a range of small-world networks with varying characteristics can be generated, and the value of $p$ can be selected based on the case. The example of an SW network shown in Figure 1(b) is generated using the WS method.
  3. 3.
    Scale-free (SF) networks (see Figure 1(c)) follow a power-law degree distribution (Figure 1(f)), indicating the presence of nodes with very high degree values, known as “hubs,” which lie in the tail of the distribution (Barabási and Bonabeau, 2003). These hubs form the core of SF networks and play a critical role in maintaining network connectivity. The removal of a hub can cause significant disconnection in the network, as a disproportionate number of paths traverse through these hubs. An example of an SF network is shown in Figure 1(c). Typically, SF networks are generated using the preferential attachment process described in (Barabási and Albert, 1999), wherein a new node preferentially connects to well-connected nodes. The probability of creating an edge to a node is proportional to its degree. The power-law degree distribution is expressed as:
    $p_{k}\propto k^{-\alpha},$ (2)
    where $\alpha$ is the distribution parameter.
2.2 Sampling algorithms
We compare the RW and the MHRW sampling algorithms for sampling the ER, SW, and SF networks, as described in Section 2.1.1. Below, we describe the two sampling algorithms:
  1. 1.
    The RW algorithm is a classic traversal-based sampling method (Göbel and Jagers, 1974). Sampling starts with selecting an initial node at random from the network. The next node is chosen uniformly from the neighbours of the initial node. These two steps are repeated until the desired sample size is achieved. The probability of choosing a neighbour $w$ of node $v$ is denoted by $P_{v,w}^{RW}$, which is given by
    $P_{(v,w)}^{RW}=\begin{cases}\frac{1}{k_{v}}&\text{ if }w\text{ is neighbour of }v,\\ 0&\text{ otherwise},\end{cases}$ (3)
    where $k_{v}$ is the degree of node $v$.
  2. 2.
    The MHRW algorithm is also a traversal-based sampling method (Hu and Lau, 2013). Sampling starts by randomly selecting a node from the network. The next node is selected from the neighbours of the initial node using the proposal-acceptance mechanism (Spencer, 2021) based on the degree of the two nodes. Suppose that the last sampled node is $v$ and the proposed next node is $w$, the probability of accepting $w$ is
    $P_{(v,w)}^{MH}=\begin{cases}\frac{1}{k_{v}}\min\Big{(}1,\frac{k_{v}}{k_{w}}\Big{)}&\text{ if }w\text{ is neighbour of }v,\\ 1-\sum_{y\neq v}P_{(v,y)}^{MH}&\text{ if }w=v,\\ 0&\text{ otherwise}.\end{cases}$ (4)
Since the RW and MHRW are Markov chain-based algorithms and share the “memoryless” property, some nodes may be revisited during the sampling process, leading to duplicates. The duplicates introduced by revisiting nodes reveal key structural features of the network. For example, they highlight community clustering, as random walks tend to revisit nodes within tightly connected groups (Rosvall and Bergstrom, 2008). Duplicates also emphasize high-degree nodes, required for immunization and intervention strategies (Maiya and Berger-Wolf, 2011). In biological and knowledge graphs, duplicates preserve important clusters and reflect the intensity of interactions, such as gene or drug-target relationships (Dempsey et al., 2012, Zhang et al., 2021).
2.3 SIR model
In Section 2.1, we introduced the structure and properties of the networks, we now demonstrate how to apply a stochastic SIR model on three networks (ER, SW, and SF) to compare disease metric estimates derived from samples obtained using the RW and MHRW algorithms.
Consider a population of size $N$, which is categorized into three compartments according to the infection status of the individuals during an epidemic. The three compartments are Susceptible ($S$), Infected ($I$), and Recovered ($R$). Let $\bm{X}(t)\equiv(S(t),I(t),R(t))$ be the state vector where $S(t),I(t),R(t)$ are the number of individuals in state $S,I,R$ respectively at time $t$. The number of individuals in each compartment of the SIR model changes due to two events: 1) a transmission event, and 2) a recovery event, which can be written, respectively as follows,
[…]
where $\beta/N$ is the transmission rate per infected individual per unit time and $\gamma$ is the recovery rate per infected individual per unit time.
The result of these two reactions/events is either the, respective, increase or decrease in the number of individuals in states $\{S,I,R\}$, which is translated as the change in the state vector $\bm{X}$. Let $\bm{v}_{i}\equiv(v_{S},v_{I},v_{R})$ be the state-change vector, where $i=1$ denotes reaction (5) and $i=2$ denotes reaction (6). The state change vector for reactions (5) and (6) are $\bm{v}_{1}=(-1,1,0)$, and $\bm{v}_{2}=(0,-1,1)$, respectively.
The time-dependent state vector $\bm{X}(t)\equiv(S(t),I(t),R(t))$ is a Markov process, i.e., the state of the system at time $t+dt$, where $dt\geq 0$, is dependent only on the state at time $t$. Thus, the time evolution of $P(\bm{X},t|\bm{X}_{0},t_{0})\equiv$ probability that $\bm{X}(t)=\bm{X}$ given that $\bm{X}(t_{0})=\bm{X}_{0}$ for any time $t\geq t_{0}$ is given below, which is the chemical master equation (CME) (Gillespie, 1992, Csefalvay, 2023).
$\frac{\partial P(\bm{X},t|\bm{X}_{0},t_{0})}{\partial t}$ $=\sum_{j=1}^{2}\Bigg{[}\underbrace{a_{j}(\bm{X}-\bm{v}_{j})P(\bm{X}-\bm{v}_{j},t|\bm{X}_{0},t_{0})}_{\text{inflow rate, probability to reach state }\bm{X}\text{ given that }\bm{X}(t)=\bm{X}-\bm{v}_{j}}$
$\quad-\underbrace{a_{j}(\bm{X})P(\bm{X},t|\bm{X}_{0},t_{0})}_{\text{outflow rate, probability to move away from state }\bm{X}\text{ given that }\bm{X}(t)=\bm{X}-\bm{v}_{j}}\Bigg{]},$ (7)
where $a_{1}(\bm{X}(t))=S(t)I(t)\beta/N$, and $a_{2}(\bm{X}(t))=I(t)\gamma$ are the propensity functions.
Due to the nonlinearity of Eq. (2.3) obtaining a closed-form solution is not possible. However, we can simulate the SIR model on networks using the Gillespie algorithm (Gillespie, 2007), which provides exact numerical realizations of the system dynamics.
2.4 Network Generation and SIR Model Simulation
Above, we have discussed the theoretical aspects of network types, network properties, sampling algorithms, and a stochastic SIR model. In this section, we provide the libraries used for generating the networks and for simulating the SIR model on networks. The pseudocode for the RW and the MHRW sampling algorithms is in Appendix A.
We use the NetworkX Python library (Hagberg et al., 2008) to generate the three networks described in Section 2.1.1. ER networks are generated using the gnp_random_graph function, SW networks with the watts_strogatz_graph function, and SF networks with the configuration_model function, where the input is a sequence of degree values drawn from a power-law distribution (using the random.zipf function from the NumPy library (Harris et al., 2020)). We generate $n=10,000$ networks for each type using the parameters listed in Table 1 to achieve the precision of $10^{-2}$ in the mean estimates of network characteristics, based on the relationship between sample size (or the number of simulations) and the standard error of the mean (Bondy and Zlot, 1976), given by
$\bar{\sigma}_{x}=\frac{\sigma}{\sqrt{n}},$ (8)
where $\bar{\sigma}_{x}$ is standard error of the mean, $\sigma$ is the population standard deviation, and $n$ is the sample size.
Network Parameter Network Type ER network SW network SF network
Number of nodes $10,000$ $10,000$ $10,000$
Average degree $5$ $5$ $5$
Edge creation Probability $(p)$ $0.5$
Power-law exponent $(\alpha)$ $3$
Table 1: Chosen parameters for generating examples of three (ER, SW, SF) networks using the NetworkX (Hagberg et al., 2008) library.
In Figure 2, examples of an ER network, a SW network, and an SF network are shown, each generated using the NetworkX library with $50$ nodes, an average degree of $5$, and other parameters as mentioned in Table 1.
We use the EoN Python library (Miller and Ting, 2020) to implement the Gillespie algorithm (Gillespie, 2007) simulating a stochastic SIR model (Section 2.3) on a network. The parameters for the SIR model include a recovery rate, $\gamma\in\{1,1/7,1/14\}$, and a transmission rate, $\beta/N$, which ranges from 00 to $1$ with a step size of $0.1$. We first consider $\gamma=1$, so that $\beta/N$ can be interpreted as a basic reproduction number for the SIR model. To understand the impact of longer recovery periods on estimates from samples, we use $\gamma=1/7$ and $1/14$, which correspond to average recovery periods for influenza (one week) (Chowell et al., 2008) and COVID-19 (two weeks) (SeyedAlinaghi et al., 2021), respectively.
After simulating the SIR model on the three networks, we obtain $100$ samples with $500$ nodes for each network type (ER, SW, and SF), using the RW and MHRW sampling algorithms (see pseudocode in Appendices 1,2). We generate $2500$ nodes per sample for each algorithm and discard the initial $2000$ nodes as burn-in nodes; to allow the sampling algorithms to reach their stationary distribution and reduce the bias from the starting node. Here, the sample size of $500$ (equivalent to $5\%$ of the population) is chosen to keep the margin of error approximately ±$\pm 5\%$ in the estimates (Conroy et al., 2016). We obtain $100$ samples from each of the $10,000$ networks in each network type, to get the precision of mean estimates up to the third decimal place (Bondy and Zlot, 1976). The code for the entire process pipeline is available on GitHub.
2.5 Hypothesis Test Statements
We perform hypothesis testing to statistically compare the estimates of disease metrics in RW and MHRW samples, relative to the underlying network (UN). The first step is to check if the data is normally distributed. We use one-sample Kolmogorov-Smirnov (KS) test (Jr, 1951) to test the following hypotheses:
Hypothesis 2.1.
Null Hypothesis (H0):
The proportion of infected nodes, for all $\beta$ values, follow a Normal distribution.
Alternative Hypothesis (H1):
The proportion of infected nodes, for all $\beta$ values, does not follow a Normal distribution.
Hypothesis 2.2.
Null Hypothesis (H0):
The number of secondary infections, for all $\beta$ values, follow a Normal distribution.
Alternative Hypothesis (H1):
The number of secondary infections, for all $\beta$ values, does not follow a Normal distribution.
Based on the results of the normality test, we then use the one-tailed Mann-Whitney U (U) test (Mann and Whitney, 1947) to test the hypotheses for comparing two samples and understand the direction of estimates from the samples relative to the UN. These hypotheses are stated as,
Hypothesis 2.3.
Null Hypothesis (H0):
  • 1.
    There is no difference between the probability distribution of the proportion of infected nodes in RW, MHRW and UN.
  • 2.
    The proportion of infected nodes in UN is greater than or equal to those in RW or MHRW.
Alternative Hypothesis (H1):
The proportion of infected nodes in the UN tends to be less than those in RW or MHRW.
Hypothesis 2.4.
Null Hypothesis (H0):
  • 1.
    There is no difference between the distribution of the proportion of infected nodes in RW or MHRW and UN
  • 2.
    The proportion of infected nodes in UN is less than or equal to those in RW or MHRW.
Alternative Hypothesis (H1):
The proportion of infected nodes in the UN tends to be greater than those in RW or MHRW.
Hypothesis 2.5.
Null Hypothesis (H0):
  • 1.
    There is no difference between the probability distribution of the number of secondary infections in RW, MHRW and UN
  • 2.
    Number of secondary infections in UN are greater than or equal to those in RW or MHRW.
Alternative Hypothesis (H1):
The number of secondary infections in the UN tends to be less than those in RW or MHRW.
Hypothesis 2.6.
Null Hypothesis (H0):
  • 1.
    There is no difference between the distribution of the number of secondary infections in RW or MHRW and UN
  • 2.
    Number of secondary infections in UN are less than or equal to those in RW or MHRW.
Alternative Hypothesis (H1):
The number of secondary infections in the UN tends to be greater than those in RW or MHRW.
3 Results
In this section, we present a comparative analysis of the RW and MHRW sampling methods—focussing on their ability to capture the degree distribution of the UN and their impact on the estimation of disease characteristics for three network types (ER, SW, and SF) relative to the UN.
Since RW and MHRW are memoryless algorithms, resulting in duplicate nodes in a sample (Hu and Lau, 2013). We compare the outcomes when duplicates are removed or retained. This analysis aims to highlight how this subtle aspect can significantly influence the accuracy of disease metric estimation across network types.
3.1 Degree Distribution
(a) Erdős-Rényi network
(b) Small-world network
(c) Scale-free network
Figure 3: The ratio of degree distributions for samples generated by the RW algorithm (red diamonds) and the MHRW algorithm (blue triangles), relative to the UN, is presented for three network types (ER, SW, SF). The UN degree distribution is based on SIR model simulations ($\gamma=1$) for $10,000$ networks of each type (parameters in Table 1) generated with NetworkX. Sample degree distribution is derived from $100$ samples (size $500$) for $10,000$ networks using RW and MHRW.
In Figure 3, we compare the degree distribution ratios of the RW and MHRW samples to the UN for the ER, SW, and SF networks. The degree distribution for the UN is based on $10,000$ networks of each type, generated using the NetworkX Python library with parameters listed in Table 1. For the samples, it is estimated from $100$ samples, each with a sample size of $500$, for $10,000$ networks of three types using the RW and MHRW algorithms. The vertical axis shows the probability ratio of a node with degree $k$ in the samples to that in the UN. The solid black line represents a ratio of $1$, so closeness to this line indicates that the sample’s degree distribution is similar to the UN. Results are consistent for samples with and without duplicate nodes (data not shown).
For the ER (Figure 3(a)) and SW (Figure 3(b)) networks, where the average degree $\geq 5$, the MHRW samples (blue triangles) align closely with the UN, while the RW samples (red diamonds) increasingly deviate as degree values increase, showing size bias. This highlights the MHRW’s accuracy in representing the UN’s degree distribution.
In Figure 3(c), both sampling methods do not perform well for the SF network. There is significant variability and both methods over-represent high degree nodes. The RW samples consistently show a higher ratio than the MHRW samples, showing more size bias. The presence of “hubs” in SF networks leads to a high variance in the degree distribution, leading to higher variability in the ratio compared to the ER and SW networks.
These results suggest that for networks with properties similar to the ER and SW, the MHRW sampling algorithm may be a more suitable choice for unbiased degree distribution. Whereas networks with SF properties require a more delegated approach, such as SLSR (Jiao, 2024), where core and peripheral nodes are sampled separately for balanced sampling.
3.2 Disease metrics
In the previous section, we compared the degree distribution of the samples generated using the RW and MHRW algorithms; we observed that the degree distribution of MHRW samples is closer to the UN than RW samples, due to bias towards high-degree nodes. Now, we compare the sampling algorithms in estimating three disease metrics: 1) proportion of infected nodes, 2) average number of secondary infections, and 3) time to infection. We simulate the SIR model on networks to estimate the disease metrics in the samples and the UN. Here, we present the results for $\gamma=1$; see C.2 for $\gamma\in\{1/7,1/14\}$.
3.2.1 Proportion of infected nodes
Panel 1: Samples with duplicate nodes.
(d) Erdős-Rényi network
(e) Small-world network
(f) Scale-free network
Panel 2: Samples without duplicate nodes.
(g) Erdős-Rényi network
(h) Small-world network
(i) Scale-free network
Figure 4: The ratio of the proportion of infected nodes during an epidemic between RW (dashed red) and MHRW (dotted blue) samples, relative to their respective UNs, is shown for (1) retained and (2) removed duplicate nodes. The light red and blue bands indicate the range of one standard deviation. The UN infection count is based on SIR model simulations ($\gamma=1$) for $10,000$ networks of each type (parameters in Table 1) generated with NetworkX (Hagberg et al., 2008). Sample infection counts are derived from $100$ samples (size $500$) for $10,000$ networks using RW and MHRW.
In Figure 4, we compare the proportion of infected nodes in the RW and MHRW samples relative to the UN as the transmission rate $\beta$ increases, with a recovery rate $\gamma=1$, for three networks (ER, SW, and SF). Panel 1 (Figure 3(d), 3(e), 3(f)) shows the results for samples with duplicate nodes, while Panel 2 (Figure 3(g), 3(h), 3(i) ) shows results for samples without duplicates. The closer the ratio is to the solid black line, the closer the sample estimates are to those of the UN.
For $\beta>0.2$, RW samples consistently overestimate the proportion of infected nodes in ER and SW networks compared to MHRW samples, regardless of the inclusion of duplicate nodes (Figure 4). In SF networks, while MHRW samples tend to overestimate the average number of infected nodes relative to RW samples, estimates from both algorithms exhibit substantial variability and fail to accurately represent the underlying network structure.
For ER networks, the MHRW samples overestimate the proportion of infected nodes when duplicate nodes are removed (Figure 3(g)) but aligns well when duplicate nodes are retained (Figure 3(d)). RW estimates remain consistent regardless of duplicate removal. Similar behaviour is observed for SW networks in Figure 3(h) and Figure 3(e).
For the SF networks, both sampling algorithms fail to sample the UN well (Figure 3(f), Figure 3(i)). In SF networks, removing duplicates increases the deviation between RW and MHRW estimates for $\beta>0.6$, with MHRW overestimating the proportion of infected nodes relative to RW. High variability in SF networks reflects challenges in capturing the heterogeneity and complex structure (e.g., hubs and peripheral nodes; Figure 3(c)).
We use hypothesis testing to statistically compare the estimates from the sampling algorithms. First, we conduct a one-sample KS test to assess whether the proportion of infected nodes follows a Normal distribution, as stated in Hypothesis (2.1), with significance level $\alpha=0.05$. The KS test statistic is summarized in B Table 5 for three data groups UN, RW, and MHRW for ER, SW, and SF networks. The KS statistic is apporximately $0.5$ across networks and data groups, with all corresponding p-values less than $0.05$. These results indicate significant deviation from normality, providing evidence to reject $H_{0}$ in Hypothesis (2.1).
Since the proportion of infected nodes does not follow a Normal distribution, we conduct a one-tailed U-test to compare the RW and the MHRW sample estimates with the UN. We test two hypotheses, one for statistically significant overestimation, Hypothesis (2.3), and another for statistically significant underestimation Hypothesis (2.4), with significance level $\alpha=0.05$. The U-test statistic provides a measure of closeness between the sample estimates and the UN, with a relatively larger value indicating greater similarity to the UN. The statistics of the U-test are summarized in Table 2 for samples with duplicate nodes and samples without duplicate nodes.
The p-values are less than $0.05$ across network types for the RW samples (with and without duplicates) and the MHRW samples without duplicates nodes, for hypothesis (2.3). This provides evidence to reject $H_{0}$ (2.3) and indicates the significant overestimation of the proportion of infected nodes. For the MHRW samples with duplicate nodes, the p-value is less than $0.05$ for the SF network, evidence to reject $H_{0}$ in Hypothesis (2.3). This indicates the significant overestimation by the MHRW samples for the SF network. On the other hand, for the MHRW samples with duplicate nodes in the ER and SW networks, the p-values are less than $0.05$ for Hypothesis (2.4) (reject $H_{0}$), indicating significant underestimation of the proportion of infected nodes.
We find that both sampling algorithms do not provide estimates statistically similar to the UN. However, for the MHRW sample estimates (with or without duplicates), the value of U-test statistics is greater than the RW sample estimates (see Table 2) for the ER and SW networks. For example, in the case of the ER networks with duplicate nodes, U-test statistics for the MHRW is $5.90\times 10^{11}$ and for the RW is $5.07\times 10^{11}$. This indicates that the MHRW sample estimates are closer to the UN compared to the RW.
For SF networks, the RW sample estimates are closer to the UN compared to the MHRW, as the U-test statistics are higher for the RW. For example, in the case of duplicate nodes, the U-test statistic is $6.1\times 10^{10}$ for RW and $2.8\times 10^{10}$ for the MHRW.
Network Sample With Duplicate Nodes Without Duplicate Nodes
RW MHRW RW MHRW
ER $5.07\times 10^{11}$ $5.90\times 10^{11}$ $5.11\times 10^{11}$ $5.38\times 10^{11}$
SW $5.8\times 10^{11}$ $5.9\times 10^{11}$ $5.82\times 10^{11}$ $5.86\times 10^{11}$
SF $6.1\times 10^{10}$ $2.8\times 10^{10}$ $5.72\times 10^{10}$ $7.83\times 10^{9}$
Table 2: U-test statistics (Hypotheses (2.3) and (2.4)) for comparing the estimates of the proportion of infected nodes in the samples generated using the RW and the MHRW sampling algorithms with the UN, with and without duplicate nodes.
Overall, while MHRW is generally more accurate for ER and SW networks, its performance degrades for SF networks, where high variability and structural heterogeneity pose challenges for both algorithms.
3.2.2 Average secondary infections
(1) Samples with duplicate nodes.
(a) Erdős-Rényi network
(b) Small-world network
(c) Scale-free network
(2) Samples without duplicate nodes.
(d) Erdős-Rényi network
(e) Small-world network
(f) Scale-free network
Figure 5: The ratio of the number of secondary infections during an epidemic between RW (dashed red) and MHRW (dotted blue) samples, relative to their respective UNs, is shown for (1) retained and (2) removed duplicate nodes. The light red and blue bands indicate the range of one standard deviation. The UN estimates are based on SIR model simulations ($\gamma=1$) for $10,000$ networks of each type (parameters in Table 1) generated with NetworkX (Hagberg et al., 2008). Sample estimates are derived from $100$ samples (size $500$) for $10,000$ networks using RW and MHRW.
In Figure 5, we compare the number of secondary infections (effective reproduction number) in the RW and MHRW samples relative to the UN as the transmission rate $\beta$ increases, with a recovery rate $\gamma=1$, for three networks (ER, SW, and SF). Panel 1 (Figure 4(a), 4(b), 4(c)) shows the results for samples with duplicate nodes, while Panel 2 (Figure 4(d), 4(e), 4(f)) shows results for samples without duplicates.
For $\beta>0.3$, RW samples (with, or without, duplicates) overestimate the number of secondary infections in ER and SW networks compared to the MHRW samples as shown in Figures 4(a), 4(b), 4(d), and 4(e).
For ER and SW networks, the removal of duplicate nodes results in overestimation by MHRW samples (Figure 4(d)) while it aligns well when duplicate nodes are retained (Figure 4(a)). However, no change is observed for RW estimates in both scenarios (with or without duplicate nodes). Similar behaviour is observed for SW networks in Figure 4(e) and Figure 4(b).
For the SF network, both sampling algorithms overestimate the number of secondary infections, with high variability (Figures 4(c) and 4(f)). However, when duplicate nodes are retained, MHRW sample estimates are closer to the UN than the RW estimates, whereas this behaviour reverses when duplicates are removed. We consistently observe that the network properties and disease metrics are highly variable and more sensitive to changes in the system parameters for the SF networks, indicating the limitation of both sampling algorithms to represent the UN well.
We conduct hypothesis testing to statistically compare the estimates from the sampling algorithms. First, we conduct the one-sample KS test to assess whether the number of secondary infections follows a Normal distribution, as stated in Hypothesis (2.2), with significance level $\alpha=0.05$. The KS test statistic is summarized in B Table 5 for three data groups that are UN, RW, and MHRW for ER, SW, and SF networks. The KS statistic ranges from $0.5$ to approximately $0.6$ across networks and data groups, with all corresponding p-values less than $0.05$. These results indicate significant deviation from normality, providing evidence to reject $H_{0}$ in Hypothesis (2.2).
Since the number of secondary infections does not follow a Normal distribution, we conduct a one-tailed U-test to compare the RW and the MHRW sample estimates with the UN. We test two hypotheses, one for statistically significant overestimation (Hypothesis (2.5)), and another for statistically significant underestimation (Hypothesis (2.6)), with significance level $\alpha=0.05$. The statistics of the U-test are summarized in Table 3 for samples with duplicate nodes, and samples without duplicate nodes.
The p-values are less than $0.05$ across network types for the RW samples (with and without duplicates) and the MHRW samples without duplicates nodes, for hypothesis (2.5). This provides evidence to reject $H_{0}$ in Hypothesis (2.5) and indicates a significant overestimation of the number of secondary infections. For the MHRW samples with duplicate nodes, p-values are less than $0.05$ for the ER and SF networks, evidence to reject $H_{0}$ in Hypothesis (2.6), implying a significant underestimation of the number of secondary infections. However, for the SW network, the p-value is greater than $0.05$ (accept $H_{0}$) for the MHRW sample estimates with duplicate nodes, for Hypotheses (2.6) and (2.5), indicating that the MHRW estimates are statistically similar to the UN.
Similar to the proportion of infected nodes, in the case of the number of secondary infections the U-test statistics (see Table 3) are higher for the MHRW sample estimates (with or without duplicates) compared to the RW for the ER and SW networks, implying that MHRW estimates are closer to the UN. For the SF networks, the RW sample estimates are closer to the UN compared to the MHRW, indicated by the higher values of U-test statistics in Table 3.
Network Sample With Duplicate Nodes Without Duplicate Nodes
RW MHRW RW MHRW
ER $2.5\times 10^{11}$ $5.6\times 10^{11}$ $2.5\times 10^{11}$ $3.33\times 10^{11}$
SW $3.99\times 10^{11}$ $6.0\times 10^{11}$ $3.91\times 10^{11}$ $4.51\times 10^{11}$
SF $8.2\times 10^{10}$ $6.6\times 10^{10}$ $8.45\times 10^{10}$ $5.99\times 10^{10}$
Table 3: U-test statistics (Hypotheses (2.5) and (2.6)) for comparing the estimates of the number of secondary infections in the samples generated using the RW and the MHRW sampling algorithms with the UN, with and without duplicate nodes.
Results for the time to infection disease metric are described in C.1. We observed that in Figure 8, estimates from both sampling algorithms (with or without duplicates) are close to the UN. Although the RW samples slightly underestimate the time to infection, whereas the MHRW samples align well with the UN.
Overall, the MHRW algorithm is more representative of the UN compared to the RW in estimating disease metrics for ER and SW networks. Thus, if the UN for real-world networks is similar to the ER and SW networks, the MHRW algorithm is recommended for sampling.
The RW samples overestimated the disease metrics regardless of whether duplicate nodes are removed or retained. The size bias in degree distribution (Section 3.1), for the RW samples, leads to overestimating the number of secondary infections for all networks, as the nodes with high degrees are more likely to become infected.
The MHRW samples align well with the UN for the ER and the SW networks, for all disease metrics, only when duplicate nodes are retained in the sample. The removal of duplicate nodes leads to an overestimation of the disease metrics (Section 3.2.1 and 3.2.2). One reason for overestimation after removing duplicates for MHRW samples is the reduction in the number of unique nodes in the MHRW samples, with a comparatively smaller reduction in the number of unique infected nodes. This may occur because the MHRW algorithm tends to get ‘stuck’ at nodes with very low connectivity due to fewer potential candidate nodes for the acceptance-rejection mechanism. Another contributing factor may be the relatively higher presence of highly connected nodes—which are more likely to become infected—in the MHRW samples compared to the UN, as observed in Figure 3.
Notably, for the SF network, both the sampling algorithms fail to obtain representative samples of the UN, resulting in high variation in estimates of the degree distribution and disease metrics. The SF network structure is complex compared to the ER and the SW networks; there is more heterogeneity in degree values. Also, the structure consists of two types of components: 1) core nodes of very high degree, and 2) peripheral nodes of low degree, which constitute the majority of the network. This diversity leads to dependency on the starting node for both sampling algorithms, leading to high variability in sample estimates.
From an application perspective, the RW sampling algorithm provides conservative estimates of disease metrics, which is beneficial for devising interventions in cases of infectious diseases with high fatality rates. For such diseases, e.g. COVID-19 and Ebola, it is preferable to contain transmission as early as possible; hence, overestimating the number of infected individuals and the effective reproduction number (i.e., the number of secondary infections) aids the implementation of timely prevention measures. However, for diseases with low fatality rates, policymakers have more time to develop policies aligned with available resources. Here, the MHRW sampling algorithm is preferable, as its estimates are closer to the UN structure. In this context, precise estimates of disease metrics is integral for efficient use of time and resources to contain the disease spread.
Furthermore, the performance of any sampling algorithm depends heavily on the complexity of the UN structure and disease parameters (such as transmission and recovery rates). For complex network structures, such as scale-free networks, sample estimates tend to be highly variable and sensitive to disease parameters, underscoring the importance of providing uncertainty intervals with estimates to policymakers and considering alternative sampling algorithms (Leskovec and Faloutsos, 2006a), depending on the availability of computational resources.
We recommend that data scientists working with policymakers examine the sensitivity of sampling algorithms to network properties and disease spread parameters before selecting an algorithm. Real-world contact networks combine features of various theoretical networks; thus, it is essential to identify which network properties should remain consistent in the sample, depending on the mechanism of disease spread. Using a combination of sampling algorithms is also advised to leverage the strengths of each method.
4 Cattle Movement Network
(a) A part of the cattle movement network.
(b) Degree distribution of cattle data.
Figure 6: Cattle movement network data.
In Section 3.2, we compared the network properties and disease metrics estimated from the RW and the MHRW sampling algorithms, on three theoretical network types. Now, we implement both sampling algorithms on movement network data. We use the cattle movement data provided by the British Cattle Movement Service (BCMS), which contains information about birth, death, and movement of the cattle.
We use cattle movement data from January 2018 to March 2018, which is recorded at day level. We consider cattle holding places (farms, marketplaces, and slaughterhouses) as nodes and an edge is created between two nodes if there is cattle movement between them. Note that movement does not occur through all the edges every day, which implies that network structure changes with time. However, for this study, we assume that all edges are active and the network is a static network.
4.1 Data Description
The data contains movement details of $1,135,502$ cattle; there are two movement records on average, per cattle. There are $46,512$ holding places, which are the nodes of the network. A static and undirected network is created with $46,512$ nodes and $159,036$ edges. Note that an edge between a pair of nodes is counted only once, in case of multiple movement records. A part of the movement network is shown in Figure 6(a); where we see a mix of large connected component, small clusters of nodes and isolated node pairs.
Panel 1: Samples with duplicate nodes.
(c) Proportion of infected nodes
(d) Number of secondary infections
(e) Time to infection (days)
Panel 2: Samples without duplicate nodes.
(f) Proportion of infected nodes
(g) Number of secondary infections
(h) Time to infection (days)
Figure 7: The ratio of disease metric estimates between sampled generated using RW and MHRW algorithm, relative to the UN for two scenarios: (1) samples with duplicate nodes, and (2) samples without duplicate nodes. The light red and blue bands indicate the range of one standard deviation. disease metric estimates are based on $500$ SIR model simulations on cattle movement network with $\gamma=1$. Sample estimates are derived from $10,000$ samples (size $2500$) for cattle movement network using RW and MHRW.
The network’s degree ranges from $1$ to $6,556$, with an average degree of $6$. Notably, $30\%$ of the nodes have a degree of $1$, while only $2.69\%$ of the nodes have a degree greater than $20$. There is only one node with $6,556$ edges, which would be a marketplace.
The degree distribution for the network is shown in Figure 6(b), up to a degree value of $20$. The degree distribution of the cattle network is fitted with an exponential distribution (Balakrishnan, 2019) with rate parameter $0.4$ and location parameter $1$ (the distribution is shifted by one unit to the right along the horizontal axis), represented by the solid black line in Figure 6(b). Other studies state that cattle movement network follows a power-law degree distribution (Christley et al., 2005, Fielding et al., 2019). One reason for this difference is the amount and period of data, we have only three months of data (from 2018) and in these studies, data for multiple years before 2018 is analysed (Duncan et al., 2022).
We observed the presence of clusters with few connections outside the cluster, which is similar to the structure of the SW networks and agrees with the results in (Fielding et al., 2019). We also observe a uniform connectivity pattern (average degree $2$) similar to the SW and ER networks if we exclude the nodes with one connection or with degree $>1000$. We observe the SF network’s trait of the presence of “hubs” in the cattle movement network, which is suggested by the presence of $0.06\%$ nodes with degree $\geq 1000$ and $30\%$ nodes with degree $1$. There are a few nodes with very high degrees, likely markets, facilitating extensive animal movement and which can potentially act as key nodes for disease spread.
4.2 Results
Due to a lack of disease data, we study a hypothetical disease spread scenario, for example, bovine tuberculosis (Bekara et al., 2014), we run $500$ simulations of the SIR model on the cattle movement network using the method and parameters from Section 2.4. The RW and the MHRW sampling algorithms are used to generate $10,000$ samples with $2500$ as sample size ($5\%$ of the network size). The number of samples, sample size, and number of SIR model simulations are chosen to adhere precision to the 2nd decimal place as in Section 3.
In Figure 7, we compare three disease metrics in RW and MHRW samples relative to the cattle network, as the transmission rate $\beta$ increases, with a recovery rate $\gamma=1$, for the cattle movement network. The three disease metrics are: 1) proportion of infected nodes, 2) number of secondary infections, and 3) time to infection. Panel 1 (Figure 6(c), 6(d), 6(e)) shows the results for samples with duplicate nodes, while Panel 2 (Figure 6(f), 6(g), 6(h)) shows those without duplicates. The closer the disease metric estimates made from the samples are to the solid horizontal black line, the closer they are to the underlying network.
We observe that the estimate of the proportion of infected nodes in an epidemic is overestimated by RW samples in both panels (with and without duplicate nodes); see Figure 6(c) and Figure 6(f). However, the estimate from MHRW samples aligns very well with cattle network when duplicate nodes are retained (Figure 6(c)), but it overestimates when duplicate nodes are removed (Figure 6(f)). In the case of samples without duplicate nodes, the estimates of the RW and the MHRW samples are the same for $\beta>0.3$.
The RW samples greatly overestimate the number of secondary infections when duplicate nodes are retained compared to the MHRW samples, which align well with the cattle network (Figure 6(d)). We observe a significant difference in the estimates when duplicate nodes are removed. The RW sample still overestimates but it is in the range of $10-20$, whereas when duplicates are retained the range is $40-130$. However, without duplicate nodes, the MHRW samples overestimate the number of secondary infections by approximately $7$ times. Also, the variability in the estimates from the MHRW samples is higher when duplicate nodes are removed compared to RW samples.
In Panel 1 of Figure 7, the estimate of time to infection is overestimated by the MHRW samples and underestimated by the RW samples. The variability in estimates is extremely high for the MHRW samples compared to the thin band for the RW sample estimates (Figure 6(e)). Interestingly, after removing the duplicate nodes, both sampling algorithms underestimate the time to get the infection, which deviates further from the cattle network estimate as the $\beta$ value increases.
Figure 7 indicates that the estimates from MHRW samples are closer to the cattle network than the RW samples. However, the variability is significantly higher for the MHRW samples for the number of secondary infections and the time to infection.
The trend of disease metric estimates on the cattle movement network is consistent with those seen in the ER and SW networks. The RW samples provide conservative estimates of disease spread, whereas the MHRW samples yield estimates closer to the underlying network when duplicate samples are retained. For highly contagious diseases like Foot and Mouth Disease, which has a mortality rate of $20\%$ in young calves (World Organisation for Animal Health, ), the RW sampling algorithm is recommended. Conversely, for diseases such as Bovine Viral Diarrhea, with a mortality rate of approximately $7\%$ (Dobos et al., 2024), the MHRW algorithm is preferable as it provides estimates more representative of the underlying network, which is particularly valuable for optimizing intervention policies.
5 Conclusions
In this study, we examined the performance of the Random Walk (RW) and Metropolis-Hastings Random Walk (MHRW) sampling algorithms for estimating disease metrics in three network types—Erdős-Rényi (ER), Small-world (SW), and Scale-free (SF). Using a stochastic SIR model, we investigated the effects of sampling biases, particularly size bias, and the inclusion or exclusion of duplicate nodes on estimates of disease spread metrics. Our analysis focused on three disease metrics: the proportion of infected individuals, the average number of secondary infections, and the time to infection.
Our findings (Section 3.2) highlight the suitability of the MHRW algorithm for networks with properties similar to the ER and SW networks. MHRW samples consistently produced estimates that closely aligned with the underlying network structure, particularly when duplicate samples were retained. In contrast, the estimates from the RW samples exhibited a significant overestimation of the proportion of infected individuals and secondary infections, while underestimating the time to infection. This makes the RW algorithm less reliable for precise estimates but useful for generating conservative predictions, which can be valuable for early interventions in high-fatality diseases like COVID-19 and Ebola.
In SF networks, the inherent heterogeneity and presence of high-degree hubs introduced substantial variability in the estimates derived from both the RW and MHRW samples. This suggests that neither method is optimal for such networks. Alternative approaches, such as separate sampling of core and peripheral nodes using algorithms like SLSR (Jiao, 2024), may be required for more accurate and balanced sampling.
The cattle movement network, which exhibits characteristics of ER, SW, and SF networks, further supported these observations. MHRW samples were more representative of the underlying network when duplicate nodes were retained, while RW samples tended to overestimate or underestimate disease metrics. For highly contagious diseases like Foot-and-Mouth Disease, the conservative estimates from RW sampling are advantageous for rapid containment measures. For less severe diseases like Bovine Viral Diarrhea, the more precise estimates from MHRW sampling better support resource-efficient intervention strategies.
Additionally, our analysis revealed that the sensitivity of both algorithms to recovery rates ($\gamma$) significantly impacts disease metric estimates (C.2). While deviations in the proportion of infected nodes diminish with lower recovery rates, variability in the estimates of secondary infections and time to infection increases with higher transmission rates ($\beta$). These findings emphasise the importance of accounting for disease-specific parameters when choosing a sampling algorithm.
To further understand the subtle differences between the RW and MHRW sampling algorithms with respect to network structure, it would be valuable to explore the impact of selecting a specific starting node for the sampling process, rather than a random selection, particularly for SF networks. As observed, the difference in disease estimates between the RW and MHRW samples decreases with a decrease in the recovery rate. It would be interesting to determine the value of the recovery rate at which this change starts, as it could guide the choice of sampling algorithm. Additionally, exploring the sensitivity of these estimates to variations in network generation parameters, such as network size, average degree, and the power-law exponent for SF networks, would provide further insight.
CRediT authorship contribution statement
Neha Bansal:
Conceptualization, Formal analysis, Investigation, Methodology, Validation, Visualization, Writing – original draft & editing. Katerina Kaouri: Conceptualization, Formal analysis, Investigation, Methodology, Project administration, Supervision, Validation, Writing – review & editing. Thomas Woolley: Conceptualization, Formal analysis, Investigation, Methodology, Project administration, Supervision, Validation, Writing – review & editing.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Funding
This work is supported by the Natural Environment, Biotechnology and Biological Sciences and Medical Research councils (NERC, BBSRC and MRC) [grant number: NE/X016714/1] as part of the One Health for One Environment: an A-Z Approach for Tackling Zoonoses (‘OneZoo’) Centre for Doctoral Training. This work is also supported by an EPSRC Impact Acceleration Account grant, under grant number EP/X525522/1.
Data and Code
Data and code are available here: https://github.com/nehabansal26/Epidemics-on-Networks
Appendix A Pseudocode for sampling algorithms
We provide the logic of the code for extracting samples using the Random Walk (RW) (Algorithm 1) and Metropolis-Hastings Random Walk (MHRW) (Algorithm 2) algorithms; refer to Section 2.2 for the mathematical formulation of the algorithms. For both algorithms, to obtain a sample from a graph, we provide three inputs: i) a graph object, which contains the index or label of nodes, and the list of connections between nodes, ii) the number of samples, and iii) the sample size, which is the number of nodes required in a sample.
Algorithm 1 Random Walk (RW) Sampling
1:Input: Graph $G$; Number of samples $W$; Sample size $S$
2:$N\leftarrow$ Number of nodes in $G$
3:$E\leftarrow$ Edges in $G$
4:Initialize transition matrix $T\leftarrow 0^{N\times N}$
5:for each edge $(i,j)$ in $E$ do
6:     if $i\neq j$ then
7:         $T[i,j]\leftarrow 1/\text{degree of node}[i]$
8:         $T[j,i]\leftarrow 1/\text{degree of node}[j]$
9:     end if
10:end for
11:$T\_non\_zero\leftarrow$ matrix of non-zero indices in $T$
12:Initialize sample matrix $sample\_array\leftarrow 0^{W\times S}$
13:$sample\_array[:,0]\leftarrow$ random sample of $W$ nodes from $G$ as seed nodes
14:for $itr=0$ to $S-1$ do
15:     for each node $i$ in $sample\_array[:,itr]$ do
16:         $next\_node\leftarrow$ randomly select a neighbour of node $i$ from $T\_non\_zero[i]$
17:         $sample\_array[i,itr+1]\leftarrow next\_node$
18:     end for
19:end for
Algorithm 2 Metropolis-Hastings Random Walk (MHRW) Sampling
1:Input: Graph $G$; Number of samples $W$; Sample size $S$
2:$N\leftarrow$ Number of nodes in $G$
3:$E\leftarrow$ Edges in $G$
4:Initialize $T\leftarrow 0^{N\times N}$
5:for each edge $(i,j)\in E$ do
6:     if $i\neq j$ then
7:         $T[i,j]\leftarrow\min\Big{(}1,\frac{\text{degree of node}[i]}{\text{degree of node}[j]}\Big{)}$
8:         $T[j,i]\leftarrow\min\Big{(}1,\frac{\text{degree of node}[j]}{\text{degree of node}[i]}\Big{)}$
9:     end if
10:end for
11:$random\_probs\leftarrow$ matrix of size $W\times S$ with random numbers from 00 to $1$
12:Initialize sample matrix $sample\_array\leftarrow 0^{W\times S}$
13:$sample\_array[:,0]\leftarrow$ random sample of $W$ nodes from $G$ as seed nodes
14:for $itr=0$ to $S-1$ do
15:     for each node $i\in sample\_array[:,itr]$ do
16:         $choices[i]\leftarrow$ list of neighbours of node $i$ where $T[i,:]>random\_probs[i,itr]$
17:         if $choices[i]$ not empty then
18:              $next\_node\leftarrow$ randomly select a node from $choices[i]$
19:         else
20:              $next\_node\leftarrow$ $sample\_array[i,itr]$
21:         end if
22:         $sample\_array[i,itr+1]\leftarrow next\_node$
23:     end for
24:end for
Appendix B Statistical test results
We discuss the results of the normality test on the data for two disease metrics: proportion of infected nodes, and number of secondary infections. We perform the normality test with the hypotheses stated in Section 2.5 on the data of disease metrics from the UN, the RW samples, and the MHRW samples. First, we assess whether the data follows a Normal distribution or not, as it influences the choice of statistical tests for comparing the disease metrics among the RW, MHRW, and UN data. If the data is normally distributed then parametric tests will be used; otherwise, non-parametric tests will be used.
We use the one-sample Kolmogorov-Smirnov (KS) test with a significance level of $\alpha=0.05$ to assess the normality hypotheses (2.1) and (2.2). The KS test statistic is the measure of the largest absolute difference between the cumulative distribution of the data (UN, RW, MHRW) and the normal distribution, which ranges from 00 to $1$. A larger KS statistic indicates a greater deviation from the normal distribution. A KS statistics value closer to $1$ with a p-value $<0.05$, suggests that the data distribution is significantly different from the normal distribution, while a value closer to 00 with a p-value $>0.05$ implies that the data is significantly similar to a normal distribution. Below are the results for the two disease metrics across three networks (ER, SW, and SF) for the data from the UN, RW, and MHRW samples.
  1. 1.
    Proportion Infected Nodes: For all three networks, the KS statistic (Table 5) values are approximately $0.5$, with p-values $<0.05$, evidence to reject the $H_{0}$ in Hypothesis (2.1) for the UN, RW and MHRW data. The KS statistic is sufficiently large, leading us to conclude that the data distributions for the proportion of infected nodes significantly deviate from a Normal distribution.
  2. 2.
    Secondary Infections: For all three networks, the KS statistic (Table 5) values are approximately $0.5$, with p-values $<0.05$, evidence to reject the $H_{0}$ in Hypothesis(2.2) for the UN, RW and MHRW data. The KS statistic is sufficiently large, leading us to conclude that the data distributions for the number of secondary infections significantly deviate from a Normal distribution.
Network UN RW MHRW
ER 0.504 0.501 0.501
SW 0.504 0.501 0.501
SF 0.504 0.502 0.507
Table 4: One-sample KS test statistics for the proportion of infected nodes (Hypothesis (2.1)).
Network UN RW MHRW
ER 0.578 0.599 0.582
SW 0.509 0.500 0.500
SF 0.509 0.500 0.500
Table 5: One-sample KS test statistics for the number of secondary infections (Hypothesis (2.2)).
Since, data does not follow a Normal distribution across networks, for both disease metrics, a non-parametric statistical test is required to compare the RW and MHRW samples with the UN. We are using the Man-Whitney U-test in this work, results for the same are discussed in Section 3.
Appendix C Disease metrics
We compare the sampling algorithms (RW and MHRW) based on the accuracy of disease metric estimates for the ER, SW and SF networks using the method described in Section 2.4. Specifically, we discuss the estimates of the time to infection for increasing $\beta$ values while keeping $\gamma=1$ in Section C.1. In Section C.2, we compare the sampling algorithms for three disease metrics: 1) proportion of infected nodes, 2) average number of secondary infections, and 3) time to infection for the ER and SW networks, with $\beta$ values from 00 to $1$ and $\gamma\in\{1/7,1/14\}$.
C.1 Time to get infection
In Figure 8, we compare the estimates of the time to infection from the RW and MHRW samples with the UN for increasing $\beta$ values and $\gamma=1$ for three networks (ER, SW, and SF). In Panel 1 (Figure 7(a), 7(b), 7(c)) results for samples with duplicate nodes and in Panel 2 (Figure 7(d),7(e),7(f)) results for samples without duplicate nodes are shown.
We observe that for ER networks and $\beta<0.2$, there is a large variation in the estimates of both the sampling algorithms in both panels. For $\beta\geq 0.2$, the MHRW sample estimates align well with the UN, while the RW sample estimates slightly underestimate the time to infection, if duplicate nodes are retained in the samples (Figure 7(a)). When duplicate nodes are removed, the MHRW sample estimates slightly underestimate the time to infection (Figure 7(d)), although the behaviour of the RW sample estimates remains consistent. For SW networks, a behaviour similar to that of the ER networks is observed. However, the large variation in the estimates for both sampling algorithms persists up to $\beta=0.5$.
Surprisingly, for SF networks, as shown in Figures 7(c) and 7(f), the behaviour of both algorithms remains consistent across both panels, i.e., estimates from both the algorithms align well with the UN and have small variation band.
Overall, both sampling algorithms are performing well across network types, particularly for the Scale-free (SF) networks, which is not the case for other two disease metrics (the proportion of infected nodes and the number of secondary infections).
Panel 1: Samples with duplicate nodes.
(a) Erdős-Rényi network
(b) Small-world network
(c) Scale-free network
Panel 2: Samples without duplicate nodes.
(d) Erdős-Rényi network
(e) Small-world network
(f) Scale-free network
Figure 8: Comparison of average time to get infected between samples generated using RW (dashed red line) and MHRW (dotted blue line), relative to the UN. (1) Duplicate nodes in the samples are retained, and (2) duplicate nodes in the samples are removed. The light red and blue bands indicate the range of one standard deviation. The UN estimates are based on SIR model simulations ($\gamma=1$) for $10,000$ networks of each type (parameters in Table 1) generated with NetworkX (Hagberg et al., 2008). Sample estimates are derived from $100$ samples (size $500$) for $10,000$ networks using RW and MHRW.
C.2 Sensitivity analysis - Recovery rate
In Section 3, we discussed the results with a varying transmission rate $\beta$ and a fixed $\gamma=1$; here we investigate the impact of varying $\gamma$ along with $\beta$ on the accuracy of disease estimates from sampling algorithms for the ER and SW networks. We compare the disease metric estimates from both sampling algorithms (RW and MHRW) with the UN, using three values of recovery rate $\gamma$ that are $1$, $1/7$, and $1/14$. We also discuss the results based on two scenarios: with duplicate nodes in the samples, and without duplicate nodes in the samples.
C.2.1 Samples with duplicate nodes
In Figure 9, we compare the average ratio of disease metric estimates between the RW and MHRW samples (with duplicate nodes) with the UN for three disease metrics: i) proportion of infected nodes, ii) number of secondary infections, and iii) time to get infected in days. The recovery rate ($\gamma$) values are $1,1/7$, and $1/14$ and transmission rate $\beta$ increases from 00 to $1$. The purple line plots with circle markers is for $\gamma=1$, the brown line plots with triangle markers is for $\gamma=1/7$, and the green line plots with square markers is for $\gamma=1/14$. For a network and a disease metric, there are two plots, one for the RW sample estimates and another for the MHRW sample estimates.
For the ER network, as shown in Figures 8(a) and 8(b), the sample estimates for the proportion of infected nodes become closer to the UN as the recovery rate decreases from $1$ to $1/14$, for both sampling algorithms. Especially, the MHRW sample estimates are well aligned with the UN for $\gamma=1/14$. For the RW sample estimates, there is a significant drop in overestimation with decrease in the recovery rate from $1$ to $1/14$. We observe similar behaviour for the SW networks, as shown in Figures 8(c) and 8(d).
For the ER network, as shown in Figures 8(e) and 8(f), the sample estimates for the number of secondary infections deviate far from the UN as the recovery rate decreases from $1$ to $1/14$, for both sampling algorithms. The MHRW sample estimates are closer to the UN compared to the RW sample estimates irrespective of the recovery rate values. We observe similar behaviour for the SW networks, as shown in Figures 8(g) and 8(h).
For the estimates of the time to infection in the ER network, as shown in Figures 8(i) and 8(j), the sample estimates deviate away from the UN for both sampling algorithms as recovery rate decreases. The overestimation increases for the MHRW samples and underestimation increases for the RW sample estimates. For the SW networks, as shown in Figures 8(k) and 8(l), underestimation increases for the RW sample estimates. Whereas the MHRW sample estimates align well with the UN with the decrease in recovery rate from $1$ to $1/14$ for $\beta<0.4$.
(1) Proportion of infected nodes.
(a) Erdős-Rényi; RW
(b) Erdős-Rényi; MHRW
(c) Small-world; RW
(d) Small-world; MHRW
(2) Number of secondary infections.
(e) Erdős-Rényi; RW
(f) Erdős-Rényi; MHRW
(g) Small-world; RW
(h) Small-world; MHRW
(3) Time to infection (days).
(i) Erdős-Rényi; RW
(j) Erdős-Rényi; MHRW
(k) Small-world; RW
(l) Small-world; MHRW
Figure 9: The average value of the ratio of three disease metrics estimates for ER and SW networks using RW and MHRW samples (with duplicate nodes) relative to the UN; for varying recovery rate $\gamma\in\{1,1/7,1/14\}$ and transmission rate $\beta$ from 00 to $1$. The UN estimates are based on SIR model simulations for $10,000$ networks of each type (parameters in Table 1) generated with NetworkX (Hagberg et al., 2008). Sample estimates are derived from $100$ samples (size $500$) for $10,000$ networks using RW and MHRW.
C.2.2 Samples without duplicate nodes
In Figure 10, we compare the average ratio of disease metric estimates between the RW and MHRW samples (without duplicate nodes) with the UN for three disease metrics: i) proportion of infected nodes, ii) number of secondary infections, and iii) time to get infected in days. The recovery rate ($\gamma$) values are $1,1/7$, and $1/14$, and transmission rate $\beta$ increases from 00 to $1$. The purple line plots with circle markers is for $\gamma=1$, the brown line plots with triangle markers is for $\gamma=1/7$, and the green line plots with square markers is for $\gamma=1/14$. For a network and a disease metric, there are two plots, one for the RW sample estimates and another for the MHRW sample estimates.
For the ER network, as shown in Figures 9(a) and 9(b), the sample estimates for the proportion of infected nodes become closer to the UN as the recovery rate decreases from $1$ to $1/14$, for both sampling algorithms. For the RW sample estimates, we observe a significant drop in overestimation with the decrease in recovery rate from $1$ to $1/14$. We observe similar behaviour for the SW networks, as shown in Figures 9(c) and 9(d).
(1) Proportion of infected nodes.
(a) Erdős-Rényi; RW
(b) Erdős-Rényi; MHRW
(c) Small-world; RW
(d) Small-world; MHRW
(2) Number of secondary infections.
(e) Erdős-Rényi; RW
(f) Erdős-Rényi; MHRW
(g) Small-world; RW
(h) Small-world; MHRW
(3) Time to infection (days).
(i) Erdős-Rényi; RW
(j) Erdős-Rényi; MHRW
(k) Small-world; RW
(l) Small-world; MHRW
Figure 10: The ratio of the average value of three disease metrics estimates for ER and SW networks using RW and MHRW samples (without duplicate nodes) relative to the UN; for varying recovery rate $\gamma\in\{1,1/7,1/14\}$ and transmission rate $\beta$ from 00 to $1$.The UN estimates are based on SIR model simulations ($\gamma=1$) for $10,000$ networks of each type (parameters in Table 1) generated with NetworkX (Hagberg et al., 2008). Sample estimates are derived from $100$ samples (size $500$) for $10,000$ networks using RW and MHRW.
For the ER network, as shown in Figures 9(e) and 9(f), the sample estimates for the number of secondary infections deviates far from the UN as the recovery rate decreases from $1$ to $1/14$, for both sampling algorithms. We observe similar behaviour for the SW networks, as shown in Figures 9(g) and 9(h).
For the estimates of the time to infection in the ER network, as shown in Figures 9(i) and 9(j), we observe that sample estimates deviate away from the UN for both sampling algorithms as the recovery rate decreases. The underestimation increases for the RW sample estimates and overestimation increases for the MHRW sample estimates. For the SW networks, as shown in Figures 9(k) and 9(l), underestimation increases for the RW sample estimates. Whereas the MHRW sample estimates align well with the UN with the decrease in recovery rate from $1$ to $1/14$ for $\beta<0.4$.
Overall, from comparing the sample estimates for three $\gamma$ values ($1$, $1/7$, and $1/14$), we observe that the estimates for both sampling algorithms are very similar when $\gamma=1/7$ and $\gamma=1/14$. However, with $\gamma=1$, there is a noticeable deviation in the sample estimates. This indicates that the sample estimates are sensitive to the recovery rate, as well as to the type of disease metric and network type.